A car starts at a rest, and accelerates to 26 meters per second in a time of 5 seconds. What is the acceleration of the car?
1
Expert's answer
2018-06-28T09:12:08-0400
We can find the acceleration of the car from the kinematic equation:
v=v_0+at,
a=(v-v_0)/t=(26 m/s-0 m/s)/(5 s)=5.2 m/s^2 .